OpenStudy (anonymous):

3, 8, ?, 21, 29, 38. What finishes the sequence?

pooja195 (pooja195):

wel what do you notice about the pattern? Going on from 21?

OpenStudy (anonymous):

21 + 8 =29, 29 + 9 = 38

pooja195 (pooja195):

yep so now go back to 3 8-3=5 3+5=8 8+6=?

OpenStudy (anonymous):

14. Can you help me with one more? @pooja195

pooja195 (pooja195):

ye

OpenStudy (anonymous):

5, 20, 10, 40, ?, 10 @pooja195

pooja195 (pooja195):

not sure... @Nnesha ?